### v1

# Forgetting: A New Mechanism Towards Better Large Language Model Fine-tuning

Abstract
Supervised fine-tuning (SFT) plays a critical role for pretrained large language models (LLMs), notably enhancing their capacity to acquire domain-specific knowledge while preserving or potentially augmenting their general-purpose capabilities. However, the efficacy of SFT hinges on data quality as well as data volume, otherwise it may result in limited performance gains or even degradation relative to the associated baselines. To mitigate such reliance, we suggest categorizing tokens within each corpus into two parts—positive and negative tokens—based on whether they are useful to improve model performance. Positive tokens can be trained in common ways, whereas negative tokens, which may lack essential semantics or be misleading, should be explicitly forgotten. Overall, the token categorization facilitate the model to learn less informative message, and the forgetting process shapes a knowledge boundary to guide the model on what information to learn more precisely. We conduct experiments on well-established benchmarks, finding that this forgetting mechanism not only improves overall model performance and also facilitate more diverse model responses.
1 Introduction
In recent years, we have witnessed emerging advancements in large language models (LLMs) [1, 2], powered by transformer-based architectures [3] with billions of parameters and extensive pre-training on trillions of tokens [4]. These models have evolved rapidly with continuous improvements in architectural design, training strategies, and scaling techniques [5]. They exhibit exceptional performance across a wide range of complex linguistic tasks, including reasoning, solving mathematics [10], summarization [11], language understanding, code generation [23, 24], question answering [15], etc.
Although powerful, LLMs still require SFT to enhance their performance in specialized tasks [16, 17, 18, 25]. SFT typically involves adapting the current LLM using conditional maximum likelihood principles on fine-tuning data comprising prompt-response pairs. However, its success heavily relies on the quality and volume of the data: Low quality can mislead the model learning [6, 8, 7, 9], introducing biases or inaccuracies that degrade performance, and small-scale datasets will hinder the model ability to generalize well [13]. On the other hand, collecting the ideal data needed for SFT can be challenging in practice. Generally speaking, task-specific data are often scarce, particularly in niche or emerging domains [13, 14], making it difficult to collect a sufficiently diverse dataset. Additionally, ensuring data quality is a non-trivial task, as it involves curating examples that are both representative and free from noise or errors. Even for humans, identifying whether the data meet high-quality standards can be difficult due to the subtleties of language and context. Consequently, the lack of high-quality, task-specific data becomes a bottleneck for SFT, limiting the potential of LLMs to excel in specialized applications.
How can we mitigate the impacts of data on fine-tuning? Data filtering [19] offers a promising solution. Specifically, it involves selecting a subset of data from the whole set that is expected to be more beneficial for the targeted LLM than the original. With proper selection rules, such as gradient behaviors [20], margins, loss, and influence [22], filtering can refine data quality effectively. However, this comes at the cost of reducing the scale of the dataset, raising open questions about the trade-off between quality and scales and its impact on the generalization of the resulting model. Existing literature has attempted to mitigate this issue by exploring data rephrasing [32, 33], while this approach heavily depends on manual efforts and/or expensive generators that are task-specific.
In this paper, we explore a new mechanism towards better LLM fine-tuning, referred to as forgetting. Following previous wisdom [12, 32, 37, 47], we begin by performing data filtering at the token level, categorizing tokens as either positive or negative based on their influence to enhancing performance. Note that token-level filtering helps preserve the data scale as much as possible, thus adopting as a default choice. Then, for positive tokens, conditional maximum likelihood is applied as usual, since our selection rules ensure that their learning will benefit the current model. Furthermore, for negative tokens, rather than simply discarding them, we propose applying forgetting (also referred to as unlearning [29, 34, 35, 36, 42, 37]) to reduce the likelihood of their generation. Compared to positive ones, negative tokens are more likely to carry uninformative or even misleading knowledge. Explicitly forgetting these tokens not only prevents the model from generating them but also helps avoid overfitting to the current corpus. Moreover, we maintain the same data scale as in conventional fine-tuning, while taking some data (tokens more accurately) as negative samples to help the model establish a clearer knowledge boundary, thereby facilitating model generalization.
Although straightforward to implement, we demonstrate the importance of forgetting in SFT for improved generalization through our extensive experiments. Specifically, we build our training corpus across 5 representative reasoning, knowledge and conversational datasets, and evaluate our forgetting mechanism alongside baseline methods on 5 diverse benchmark datasets, incorporating various LLMs as base models. For example, as shown in Table 3 in Section 5, using LLaMA3.2-1B as the base model, our approach achieved a 2.51% improvement over that without forgetting and a 4.49% improvement over the fine-tuned model on full tokens. Similarly, with LLaMA3.2-3B, we obtained a 3.4% improvement over that without forgetting and 5.28% over fine-tuned model on full tokens. Additionally, with LLaMA3.1-8B, our approach resulted in a 4.21% improvement over the no forgetting approach, and a 8.25% improvement over the fine-tuned model on full tokens. Furthermore, we demonstrate our general effectiveness and robustness across other SFT setups in Section 5.
Connection with broader literature. The mechanism of forgetting is closely connected to preference optimization (PO) [70]. Recalling that, many representative PO methods, such as direct preference optimization (DPO) [70] and proximal policy optimization (PPO) [69], can broadly be reviewed as combining the objectives of learning and forgetting. They aim to increase the likelihood of generating preferred corpora while reducing that of the dispreferred one. However, these methods are derived from the original PO objectives, which are inherently tied to problem setups and rely on manual labeling or reward models for preference annotation. In contrast, we focus on the SFT problems, where the forgetting mechanism acts as an enhancement strategy rather than a indispensable component of the problem formulation. Our method is inspired by PO but more focuses on the mechanism of forgetting as an integral component within learning. This approach helps mitigate the negative effects of low-quality data meanwhile enhancing generalization and diversity. In the long term, we aim to bridge the methodological gap between SFT and PO, striving for a more unified and flexible framework for adapting LLMs.
2 Related works
2.1 Data selection for SFT
SFT is a well-known fine-tuning technique that maximizes the likelihood of generating target tokens under the assumption that all tokens are informative. However, data quality has emerged as a critical bottleneck for this approach [50], with errors arising from various sources including human annotators, tool annotators, LLM hallucinations, and data processing inconsistencies [50].
LIMA [26], hypothesized that LLMs primarily learn the style of dataset responses, rather than updating their pre-trained knowledge toward specialized tasks, by showing that fine-tuning on a 10k carefully curated dataset, they can obtain better performance than fine-tuning on a larger dataset.
To address quality challenges, researchers have investigated the advantages of data quality over quantity, proposing selection algorithms based on quality and diversity metrics to filter misleading samples and improve instruction-following capabilities [27, 21, 28, 31, 46]. While effective at improving performance, these approaches suffer from a fundamental limitation: they operate at the sample level, discarding entire examples and thus reducing the overall data scale available for training. This creates an inevitable trade-off between quality and quantity that remains unresolved.
Several data quality metrics have been introduced, such as gradient matching [49], human feedback [55] and influence function scores [46]. Moreover, [52] demonstrated that naturally higher influence scores for certain tasks can introduce bias in data selection, and proposed normalizing influence scores across different tasks before iteratively selecting samples for underrepresented skills. In [50], authors propose a two-stage noise-robust framework that performs noise detection using multiple expert systems and then relabels the downstream task data by finding similar examples from the clean set to provide context. In another approach, researchers showed that selecting training samples aligned with the model’s existing knowledge can improve performance by generating multiple instruction-response pairs and choosing those with the highest probability according to the target model [51].
Recent studies have explored various high-quality data selection algorithms for LLM fine-tuning, yet they predominantly overlook a crucial insight: even in noisy samples, some tokens still contain valuable information. By discarding entire samples, these methods inadvertently remove useful training signals. Furthermore, these approaches fail to utilize the rejected data as a learning signal.
2.2 LLM unlearning and PO
Several approaches have been proposed to remove specific information from LLM without complete retraining them from scratch, including data replacement and relabeling strategies [32, 33], and knowledge editing techniques by predicting targeted parameter updates to change specific facts while preserving other knowledge [34]. Gradient ascent (GA) based methods are usually used for their simplicity, which maximize the negative log-likelihood of specific token sequences[35, 36, 42, 44, 45, 30]. However, some of them lead to degradation in LLM’s outputs globally and damage the overall integrity of LLMs when removing targeted knowledge [30, 40, 41, 39, 38]—called excessive unlearning, which some regularization techniques such as minimizing the KL-Div between the output distributions of the pre-trained and fine-tuned models [43] is proposed to maintain performance on retain dataset. This introduce additional computational overhead and hyperparameter sensitivity. Researchers in [37] introduced WGA, which applies confidence-based weights to mitigate the excessive unlearning on a controlled forgetting manner.
In the PO field, DPO has emerged as an alternative to PPO-based alignment methods. However, PPO has been successful for its sample efficiency compared to earlier policy gradient methods, it still suffers from explicitly modeling a reward model, and complex hyperparameter tuning [69]. To address these challenges and making it more robust and less computationally expensive, DPO formulates the alignment objective into a maximum likelihood formulation on a preference-paired data, trying to make preferred responses more likely and dispreferred responses less likely. There are extensive studies to address the limitations of DPO [71, 72, 73, 74, 75, 76], a new approach for preference-based unlearning was proposed by [36], which defines the forget set as the dispreferred responses, and the preferred response contains the refusal responses like "I do not know the answer". Inspired by this research, [39] proposed a new variant of DPO, called negative preference optimization (NPO) that uses only negative responses, disregarding the positive ones. In the [37] further proposed Token-level NPO (TNPO) and Weighted TNPO (WTNPO), applying unlearning at the individual token level for more precise control over knowledge removal, yet these methods were developed specifically for targeted forgetting rather than as a complement to learning during SFT.
3 Preliminary
In this section, we present the foundational background essential to our work. We start by introducing SFT for autoregressive language modeling, followed by discussing the data quality issues within SFT.
3.1 SFT
Autoregressive language modeling, known as sequential prediction of outputs conditioned on previous context, plays a dominant role in contemporary LLMs. After pre-training, SFT is typically adopted to further improve LLMs for specific tasks by optimizing on task-specific instruction-response pairs. Specifically, representing a training corpus as , including sequence sample pairs, each pair containing as an input prompt and as a completion response. Each prompt is denoted as with indicating the sequence length of the -th prompt. Similarly, each -th completion response with sequence length of is denoted as . In an autoregressive manner, the model learns to estimate the probability distribution for each token in the response, conditioned on the entire prompt and all preceding generated tokens in the response , where denotes the model parameters.
The standard cross-entropy objective is typically adopted for SFT, following the formulation of
where the index set is defined as:
and the per-token loss function is defined as:
3.2 Data Quality of SFT
LLMs acquire diverse capabilities and knowledge representations through pretraining on extensive corpora. However, for utilizing them in specialized tasks, techniques such as SFT play a remarkable role in enhancing their performance by fine-tuning the LLM on the training corpus without any selection or discarding on the dataset’s components [48, 19].
However, collecting high-quality data, representing the required specific knowledge, is crucial to prevent inaccuracies and effectively align the LLM [19]. High-quality data collection can be challenging in practice due to several factors. Generally, task-specific data are often scarce, particularly in emerging domains. In addition, datasets are collected from various resources, often leading to inconsistent linguistic styles and quality, and errors due to the use of annotator tools, human manual annotating [50]. Therefore, each of them can contribute noisy and misleading tokens into the dataset thus jeopardizing the optimization process, leading to poor generalization.
To mitigate the impacts of low-quality and misleading data/tokens, existing methods proposed various data selection methods to maintain beneficial and high-quality data for fine-tuning [19]. More specifically, existing methods address data filtering at the data level; however, token-level filtering seems to preserve dataset scale and fine-grained information much more.
Although progress has been made in previous studies, they discard the low-quality data during fine-tuning, which significantly reduces the original dataset scale and potentially limits the model generalization. This remains an open question: how to leverage the full training dataset at its original scale while improving model performance? Specifically, is it possible to not only learn from high-quality samples but also utilize misleading data/tokens to establish clearer knowledge boundaries without overfitting to noise? Such an approach could lead to improvements in model generalization while maintaining the comprehensive scope of the original dataset.
4 Method
SFT is a well-established approach for aligning extensively knowledge-augmented pretrained LLMs with specialized tasks. As discussed in Section 3.2, practical datasets make it challenging for SFT to achieve high performance, as their collection process leads to a noisy dataset that jeopardizes the optimization process through misleading gradients. While many studies have attempted to address this issue by selecting high-quality subsets from SFT training data, these approaches sacrifice dataset scale instead of taking advantage from noisy tokens. This remained an open challenge to mitigate the effect of misleading tokens in the dataset, while preserving its scale. In this study, we propose a new approach for better LLM supervised fine-tuning, based on forgetting mechanism. Unlike traditional data selection approaches that treat all tokens uniformly and discard low-quality data, our method explicitly distinguishes between informative (positive) and uninformative or misleading (negative) tokens at a granular level. This token level approach preserves training data scale, while utilizing the tokens’ training signals more effectively.
Specifically, actively forgetting negative tokens, rather than merely ignoring them, can significantly improve model performance by aligning better with target data, freeing up model capacity from undesired patterns, and preventing overfitting to noisy patterns. This insight particularly valuable when working with practical datasets that inevitably include noisy tokens that should be forgotten to preserve the model’s general capabilities. The overall pipeline is outlined in Algorithm 1. In the following parts, we introduce the components of our pipeline, including the data preprocessing and training objective function.
4.1 Token quality assessment
To quantify token quality, we leverage the concept of influence functions [47], between the base and reference models. Given a base model with parameters and a reference model with parameters (introduced in Section 5.1.2), we define the cross-model influence for token as follows.
The intuition is that tokens that become more predictable after initial training (resulting in loss reduction) represent patterns that the model has successfully learned and are likely to be informative.
The token quality score formulation is as follows:
A positive quality score indicates that the token became more predictable on the reference model (lower loss in than in ), indicating that it represents a generalizable pattern. In contrast, a negative score suggests that the token might represent noise or misleading information.
4.2 Token selection
As a preprocessing step, we partition the tokens into positive and negative sets based on the quality scores. We first compute quality scores for all tokens in the training corpus, then sort them in descending order to form the set . Given a proportion hyperparameter , we partition the tokens as follows:
where denotes the -th percentile threshold in . The top proportion of tokens are considered as positive tokens form the set, while the remaining tokens form the negative set . In practice, we found that setting in the range of to achieves best results in our experiments. Furthermore, our experiments reveal that partitioning tokens by a zero threshold score (i.e. as positive tokens) negatively affects performance. This challenges the intuition that tokens with higher confidence improvement are informative and beneficial, while the others are harmful, introducing an open challenge for proposing more robust methods to identify high-quality tokens.
4.3 Training objective
While standard SFT algorithms maximize the likelihood over all tokens uniformly (potentially reinforcing noisy patterns that mislead optimization) and data selection methods discard the distinguished noisy data before training, our approach maintains the benefits of full-scale training while addressing quality concerns, which enables the model to establish clearer knowledge boundaries, by minimizing the likelihood of generating the noisy tokens and freeing model capacity from misleading noisy patterns. As mentioned in the Section 2.2, unlearning techniques proven to be effective to mitigate the influence of undesirable data while preserving the model utility. In our context, rather than forgetting some specified knowledge(e.g., copyrighted content), we forget misleading tokens through GA, effectively utilizing both positive and negative tokens. This approach enhances the model generalization while maintaining the original data scale with no information loss. We propose a training objective for our selective learning and forgetting as follows.
where the first term represents the average weighted loss over positive tokens, and the second term represents the average weighted loss over negative tokens. We use as an adaptive coefficient that scales linearly with training progress, ensuring an effective balancing of positive and negative gradients through the optimization process.
In this training objective, optimization initially shares goals with generalization, but their objectives later diverge. The forgetting mechanism acts as a regularization technique that pulls optimization back for generalization when their goals conflict. By using the adaptive balancing coefficient, this enables to better capture the underlying preferred data distribution rather than overfitting to the noise or merely following the pattern of low-scale high-quality data.
5 Experiments
5.1 Experimental setups
5.1.1 Datasets
Training data.
We constructed our training corpus by randomly sampling from five datasets, Flan_v2 [53], Dolly [54], Open Assistant 1 [55], Stanford Alpaca [56] and WizardLM [57]. Please refer to Appendix A in the Supplementary Material for more dataset details. The dataset distribution presented in detail in Table 1. This corpus provides a comprehensive coverage of domains and response styles, thereby enhancing the model’s generalization capabilities [66].
Evaluation benchmarks.
For the evaluation part, we have performed comprehensive evaluations on five diverse benchmark datasets. They are TruthfulQA [58] to evaluate the ability of LLM in providing truthful and accurate information, BoolQ [59] a binray question-answering dataset and evaluates LLM’s ability in making precise boolean judgements, LogiQA [60] focused on logical reasoning, TydiQA [63] to evaluate the LLM on multilingual question-answering and ASDiv [64] to evaluate the LLM on math word problems. The benchmarks’ attributes are presented in Table 2. The evaluation is processed on all benchmark samples, by using the lm-eval-hareness111https://github.com/EleutherAI/lm-evaluation-harness repository.
| Dataset | 50k Sample | 10k Sample | ||
| Samples | Percentage | Samples | Percentage | |
| Dolly | 2,617 | 5.23% | 503 | 5.03% |
| Flan_v2 | 17,803 | 35.61% | 3,593 | 35.93% |
| Open Assistant 1 | 5,960 | 11.92% | 1,135 | 11.35% |
| Stanford Alpaca | 9,276 | 18.55% | 1,834 | 18.34% |
| WizardLM | 14,344 | 28.69% | 2,935 | 29.35% |
5.1.2 Models
Base models.
In this paper, we choose 3 open-source LLMs including LLaMA-3.2-1B, LLaMA-3.2-3B and LLaMA-3.1-8B [65] in diverse complexity as our base models for fine-tuning.
Reference models.
The reference models are obtained by fine-tuning the base models on a subset where is the training corpus and is a combination of training datasets. The fine-tuned LLM will be used for calculating the influence scores. We also investigate the robustness of our approach when the reference dataset contains duplicate samples (see Appendix B.2).
Baselines.
In this study, our baselines include the base model, the supervised fine-tuned version of the base model on the whole training dataset with full tokens, and the fine-tuned version of the base model on the preprocessed training dataset including only the top k% clean tokens.
5.1.3 Training configurations
For the reported results in Table 3, we employed model-specific hyperparameter pairs as follows: for LLaMA-3.2-1B and for both LLaMA-3.2-3B and LLaMA-3.1-8B, for our adaptive balancing coefficient . These values were determined through ablation studies optimizing for performance across our benchmark tasks. For fine-tuning the LLMs, we used LoRA [67] for its memory efficiency and stability during training. We set rank-size of 64, the scaling factor of 16 and dropout 0.1 for LoRA. We used the AdamW optimizer [68], with the overall batch size equal to 24 and the fine-tuning process is performed for 1 epoch with a learning rate and a linear learning rate scheduler with 0.03 warm-up ratio. Moreover, we conducted our experiments on 4 NVIDIA L40S-48GB GPUs with Intel Xeon 6338 CPUs, running on Ubuntu 20.04.6 LTS. The systems utilize Transformers version 4.51.3 and CUDA version 12.5. Training time for 1B, 3B and 8B models approximately takes 2, 3, and 5 hours, respectively.
| Dataset | Focus Area | Data Size | Question Length |
| TruthfulQA | Truthfulness | 817 | Medium |
| BoolQ | Boolean QA | 15,942 | Short |
| LogiQA | Logical reasoning | 8,678 | Medium |
| TydiQA | Multilingual QA | 204k | Varied |
| ASDiv | Math Word Problem Solving | 2,305 | Varied |
| Method | TruthfulQA | BoolQ | LogiQA | TydiQA | ASDiV | AVG |
| Base model: LLaMA-3.2-1B | ||||||
| Base | 37.83±0 | 63.80±0 | 22.17±0 | 14.36±0 | 0±0 | 27.63±0 |
| Full Tokens | 38.74±0.39 | 59.84±0.94 | 24.60±0.25 | 28.10±0.46 | 0.55±0.48 | 30.37±0.39 |
| Ignoring (seq-level) | 39.56±0.57 | 61.47±0.06 | 24.03±0.25 | 27.90±0.34 | 1.46±0.15 | 30.88±0.28 |
| Forgetting (seq-level) | 38.93±0.08 | 63.13±0.46 | 24.80±0.12 | 28.75±0.23 | 2.50±0.04 | 31.62±0.10 |
| Ignoring (token-level) | 42.40±0.13 | 60.21±1.66 | 24.34±0.31 | 33.87±0.64 | 0.91±0.2 | 32.35±0.46 |
| Forgetting (token-level) | 44.83±0.45 | 65.39±0.39 | 25.60±0.48 | 36.21±0.77 | 2.28±0.04 | 34.86±0.22 |
| Base model: LLaMA-3.2-3B | ||||||
| Base | 39.45±0 | 73.04±0 | 22.17±0 | 21.12±0 | 31.24±0 | 37.40±0 |
| Full Tokens | 42.95±0.47 | 72.54±0.59 | 25.51±0.21 | 44.04±0.27 | 49.46±0.14 | 46.90±0.16 |
| Ignoring (seq-level) | 40.58±0.54 | 72.93±0.28 | 24.36±0.47 | 44.82±1.03 | 49.11±0.29 | 46.36±0.41 |
| Forgetting (seq-level) | 40.95±0.30 | 77.80±0.38 | 25.27±0.23 | 47.52±0.45 | 49.83±0.93 | 48.27±0.06 |
| Ignoring (token-level) | 47.23±0.86 | 75.40±0.37 | 25.12±0.31 | 47.63±0.42 | 48.51±0.74 | 48.78±0.19 |
| Forgetting (token-level) | 50.32±0.96 | 76.66±0.07 | 27.09±0.37 | 56.36±0.06 | 50.47±0.3 | 52.18±0.12 |
| Base model: LLaMA-3.1-8B | ||||||
| Base | 45.08±0 | 82.15±0 | 26.51±0 | 46.67±0 | 12.93±0 | 42.67±0 |
| Full Tokens | 44.51±0.48 | 81.44±0.47 | 25.68±0.14 | 52.03±0.18 | 51.46±0.42 | 51.02±0.11 |
| Ignoring (seq-level) | 47.05±0.21 | 85.17±0.45 | 24.64±0.18 | 52.34±0.08 | 51.62±0.28 | 52.16±0.24 |
| Forgetting (seq-level) | 47.83±0.09 | 85.56±0.18 | 24.85±0.37 | 57.56±0.33 | 57.76±0.18 | 54.71±0.10 |
| Ignoring (token-level) | 52.38±0.22 | 82.76±0.07 | 25.53±0.11 | 56.66±0.06 | 57.95±0.35 | 55.06±0.16 |
| Forgetting (token-level) | 58.39±0.65 | 83.14±0.15 | 31.15±0.86 | 66.21±0.23 | 57.48±0.12 | 59.27±0.35 |
5.2 Empirical Results
We conducted comprehensive experiments to evaluate our forgetting approach against all baselines. Remarkably, our method outperformed all baselines in average performance. The forgetting method achieved superior results with in the range of 70% to 80%, while the ignoring has its best-case performance with in the range of 50% to 60% across all benchmarks. We demonstrate the results of our experiments utilizing three different variants of LLaMA in Table 3, comparing the method in their best-case performance, specifically, setting for our forgetting approach and for the ignoring approach. Notably, compared to the standard SFT our method has achieved an average performance improvement of on the 1B model, on the 3B model and on the 8B model. Furthermore, compared to ignoring baseline, our method has achieved performance improvement of on the 1B model, on the 3B model and on the 8B model. Additional experiments with LLaMA-2-13B confirms these forgetting mechanism’s generalization capability in larger scales, with detailed results provided in Appendix B.1.
Token-level vs. sequence-level granularity. A key design choice in our approach is operating at the token level rather than the sequence level. This granular approach is motivated by the observation that individual sequences often contain a mixture of both informative and misleading tokens. Sequence-level selection would classify entire sequences as either positive or negative, potentially discarding valuable tokens within otherwise noisy sequences, or conversely, retaining harmful tokens within generally useful sequences. Token-level selection allows us to preserve beneficial information while selectively forgetting problematic content, maximizing the utility of our training data. The Table 3 shows a comparison of the different approaches.
Table 3 shows that token-level approaches consistently outperform their sequence-level counterparts across all model sizes. For example, with LLaMA-3.2-3B, token-level forgetting achieves 52.18% average performance compared to 48.27% for sequence-level forgetting. This superiority stems from token-level selection’s ability to preserve useful information even in partially noisy sequences, while sequence-level selection discards entire sequences that may contain valuable tokens alongside problematic ones.
5.3 Ablation study
Impact of . Our empirical evidence indicates that the forgetting approach demonstrates superior generalization capability when has a higher value, partitioning a larger subset of tokens as positive tokens and treating all remaining tokens as negative tokens (forget rate of ). However, forgetting only a subset of the remaining tokens and discarding the others leads to suboptimal performance, indicating the effectiveness of forgetting all the tokens as negative tokens. Figure 1(b) illustrates the average performance for different forget rates. Moreover, the choice of the hyperparameter , directly affects the noise distribution in positive and negative sets. Higher value of can introduce noisy tokens to the positive set, while lower value of can add informative tokens to the negative set. Figure 1(a) shows the comparison between different values of for the forgetting and ignoring approaches. The average performance of the forgetting method has significantly decreased for the lowest value , due to the higher proportion of informative tokens in the negative set.
Impact of (step). As explained in Section 4, effectively balancing the training and forgetting gradients is crucial for optimization stability. As related studies typically use a constant coefficient in the range (0,1) to reduce the learning rate of forgetting gradients. However, through empirical investigation, we observed that as training iterations progress, the learning rate reduction leads to the vanishing of the forgetting gradients. Thus, we used an adaptive function , as a coefficient on forgetting loss term of our dual objective function, not only to balance the learning and forgetting gradients, but also to efficiently preserve the effects of forgetting gradients during fine-tuning. According to the dual objective function formula, ignoring approach is equivalent to forgetting with a balancing coefficient of zero.
In a comparison of balancing coefficient strategies, we evaluated three approaches: static approaches with constant values zero (ignoring) and 0.0001 (optimal value for static strategy), and a dynamic approach using the linear function with and . The corresponding average improvements are 48.78%, 49.59%, and 52.18%, respectively. These results demonstrate that adaptive adjustment via linear function significantly outperforms static coefficient assignment, highlighting the critical role of selecting an appropriate balancing coefficient strategy. By incorporating , the forgetting learning rate decreases more gradually with a shallower slope. We investigated the impact of the adaptive parameter through a series of experiments.
Hyperparameter sensitivity analysis. To evaluate the robustness of our approach to hyperparameter choices, we conducted extensive experiments varying the key parameters and while keeping fixed. As shown in Figure 1(a), our method demonstrates impressive robustness to values across a wide range. For practical selection of , users can use the ratio of tokens with positive influence scores as an initial estimate—in our experiments, this ratio was 0.67, leading us to select as optimal. Comprehensive results across different combinations of and values using LLaMA-3.2-3B are presented in Appendix B.3.
Impact of forgetting. As demonstrated in previous sections, the forgetting mechanism significantly improves the performance of fine-tuning with respect to that without forgetting and standard SFT. Specifically, when comparing the forgetting and ignoring approaches with the same selection ratio ( = 0.7), the forgetting method achieves an accuracy of 52.18%, outperforming the ignoring approach (48.39%). This performance gap indicates that the negative tokens set has a high noise ratio, reinforcing the impact of forgetting misleading tokens, leading to higher performance.
6 Limitations
Despite our method’s improvements, some limitations remain. The approach is sensitive to dataset size and noise ratio, leading to performance degradation for smaller negative token sets. Moreover, it requires careful balancing between learning and forgetting gradients. Although it includes small number of hyperparameters, results heavily depend on hyperparameter selection.
7 Conclusion
This paper aims to reduce the reliance of LLM fine-tuning on data quality, an important and on-going topic that has been receiving increasing attentions these days. Unlike previous works that primarily focus on improving data selection, we suggest that exploring new learning paradigms is equally crucial. Specifically, we propose a novel fine-tuning mechanism named forgetting, which explicitly enables the model to forget misleading message carried by those filtered-out tokens. It mitigates the negative impact of noisy or misleading data while preserving the dataset scale, encouraging the model to form clearer knowledge boundaries and improving generalization and overall performance. In the future, we will explore more formal and rigorous ways to defining and enhancing data quality, as well as extend the forgetting mechanism to other related areas within LLMs, such as pre-training, preference optimization, and inference.
Appendix A Datasets details
Table 4 provides comprehensive information about the datasets used to create training corpus, including their quality assessment, size, total length of samples, and source.
| Dataset | Data Quality | Size | Length | Resource |
| Dolly | High | 15.01k | Varied | Human-annotated |
| Flan_v2 | High | 100k | Varied | Human-annotated |
| Open Assistant 1 | Moderate | 33.92k | Varied | Human-annotated |
| Stanford Alpaca | High | 52k | Varied | LLM-generated |
| WizardLM | High | 100k | Longer | LLM-generated |
Appendix B Additional experimental results
B.1 LLaMA-2-13B results
To further validate the robustness and scalability of our forgetting mechanism, we conducted additional experiments using LLaMA-2-13B as the base model. These results provide additional evidence that our approach consistently improves performance across different model architectures and scales, extending beyond the LLaMA-3.x series reported in the main paper.
The results in Table 5 demonstrate that our forgetting method maintains its effectiveness with larger models, achieving a 6.16% improvement over standard SFT and a 4.16% improvement over the ignoring baseline. This consistency across model scales (from 1B to 13B parameters) reinforces the generalizability of our approach and suggests that the forgetting mechanism provides fundamental benefits for supervised fine-tuning regardless of model size or architecture.
| Method | Dataset | |||||
| TruthfulQA | BoolQ | LogiQA | TydiQA | ASDiV | AVG | |
| Base model: LLaMA-2-13B | ||||||
| Base | 36.73 | 80.67 | 26.05 | 34.27 | 0.35 | 35.61 |
| Full Tokens (standard SFT) | 42.65 | 82.24 | 27.44 | 36.77 | 8.76 | 39.57 |
| Ignoring | 43.01 | 84.50 | 27.29 | 38.39 | 15.34 | 41.71 |
| Forgetting (Ours) | 52.82 | 84.13 | 27.95 | 48.71 | 17.80 | 46.28 |
B.2 Impact of Reference Dataset Duplicates
We conducted additional experiments to investigate the robustness of our approach when the reference dataset contains duplicate samples. However our pipeline’s preprocessing step removes duplicate samples from the both training and references datasets, this analysis is important for understanding how data quality in the reference model training affects the overall forgetting mechanism performance.
Table 6 shows results using LLaMA-3.2-3B when the reference dataset includes duplicate samples. Interestingly, our forgetting method remains effective even under these suboptimal reference conditions, achieving a 4.93% improvement over standard SFT and a 2.05% improvement over the ignoring baseline. This demonstrates the robustness of our influence-based token quality assessment even when the reference model is trained on imperfect data, suggesting that our approach can handle practical scenarios where perfect data curation is not feasible.
| Method | Dataset | |||||
| TruthfulQA | BoolQ | LogiQA | TydiQA | ASDiV | AVG | |
| Base model: LLaMA-3.2-3B (Reference with Duplicates) | ||||||
| Base | 39.45±0 | 73.04±0 | 22.17±0 | 21.12±0 | 31.24±0 | 37.40±0 |
| Full Tokens (standard SFT) | 42.95±0.47 | 72.54±0.59 | 25.51±0.21 | 44.04±0.27 | 49.46±0.14 | 46.90±0.16 |
| Ignoring | 49.91±0.39 | 75.60±0.86 | 24.99±0.35 | 48.61±0.20 | 49.81±0.01 | 49.78±0.22 |
| Forgetting (Ours) | 51.09±0.54 | 77.00±0.09 | 26.57±0.08 | 54.88±0.29 | 49.60±0.14 | 51.83±0.11 |
B.3 Hyperparameter sensitivity analysis
Table 7 presents comprehensive results across different combinations of and values using LLaMA-3.2-3B. The results demonstrate remarkable stability, with performance variations remaining small across different hyperparameter settings (standard deviation < 0.5% across configurations). This robustness ensures that our method maintains superiority over baselines without requiring extensive hyperparameter tuning. The stability is partly attributed to the inherent robustness of large language models and their extensive pre-trained knowledge, which provides a strong foundation that is resilient to moderate changes in fine-tuning parameters.
| TruthfulQA | BoolQ | LogiQA | TydiQA | ASDiV | AVG | ||
| 0.00001 | 0.45 | 52.75 | 74.38 | 25.89 | 54.27 | 48.10 | 51.08 |
| 0.00001 | 0.35 | 51.55 | 75.11 | 26.15 | 56.74 | 48.42 | 51.59 |
| 0.00001 | 0.25 | 50.93 | 76.58 | 25.99 | 56.13 | 50.26 | 51.98 |
| 0.00001 | 0.15 | 50.17 | 75.45 | 26.19 | 54.37 | 50.67 | 51.37 |
| 0.0001 | 0.45 | 50.90 | 77.56 | 25.83 | 54.33 | 48.90 | 51.50 |
| 0.0001 | 0.35 | 51.20 | 75.67 | 26.65 | 57.21 | 48.78 | 51.90 |
| 0.0001 | 0.25 | 50.32 | 76.64 | 27.09 | 56.36 | 50.47 | 52.18 |
| 0.0001 | 0.15 | 50.09 | 74.79 | 25.27 | 55.21 | 51.82 | 51.44 |
| 0.001 | 0.15 | 49.05 | 76.03 | 26.36 | 54.85 | 51.49 | 51.56 |
| 0.001 | 0.25 | 48.96 | 76.50 | 28.68 | 56.35 | 49.66 | 52.03 |
| 0.001 | 0.35 | 51.25 | 74.41 | 26.51 | 56.58 | 50.05 | 51.76 |
| 0.001 | 0.45 | 50.69 | 74.50 | 25.98 | 56.97 | 48.46 | 51.32 |
| 0.01 | 0.15 | 50.46 | 75.24 | 26.12 | 54.17 | 50.95 | 51.39 |
| 0.01 | 0.25 | 51.02 | 76.28 | 27.75 | 55.48 | 49.93 | 52.09 |
| 0.01 | 0.35 | 52.78 | 74.44 | 25.58 | 55.92 | 48.30 | 51.40 |
| 0.01 | 0.45 | 50.09 | 74.87 | 27.60 | 54.69 | 48.68 | 51.19 |
Appendix C A formal motivation for forgetting
This section provides theoretical foundations showing that our approach has principled convergence guarantees and optimal generalization properties. We present two complementary theoretical perspectives: (1) convergence analysis showing our dual objective converges to good solutions, and (2) domain adaptation theory explaining why token-level forgetting improves generalization across distributions.
C.1 Pareto Convergence
C.1.1 Assumptions
Assumption 1 (Regularity Conditions).
We require the following assumptions
1.
Smoothness: Each is -Lipschitz in and has -Lipschitz gradients
2.
Bounded Domain: We optimize over a bounded domain with diameter
3.
Bounded Gradients: for all
4.
Scheduling Bounds: for all
5.
Token Quality: The partition into and is fixed throughout training
Remark 2.
These assumptions are more realistic for neural network training than assuming convexity. The bounded domain can be enforced via projection or weight clipping.
Why these assumptions matter: These conditions might seem abstract, but they capture realistic training scenarios. Assumption 1-2 say our loss functions are smooth (gradients don’t change too drastically), which holds for standard neural networks. Assumptions 3-4 ensure gradients and parameters stay bounded during training, which can be enforced through gradient clipping and weight decay. Assumption 5 simply says we keep our token categorization fixed during training (rather than re-computing influence scores every step).
C.1.2 Convergence Analysis
The key question: Does our forgetting-based training actually converge to good solutions? Standard optimization theory doesn’t directly apply because our objective changes over time due to the adaptive schedule. We prove convergence using dynamic regret analysis, which measures how well we perform compared to a sequence of time-varying optimal solutions.
Our main theoretical result establishes convergence in terms of dynamic regret against time-varying optimal solutions.
Theorem 3 (Dynamic Regret Bound).
Under Assumption 1, running gradient descent with learning rate on the sequence of objectives produces iterates satisfying:
where .
Proof.
We decompose the dynamic regret as:
Static Regret Bound
Standard online gradient descent analysis gives:
Since , we get:
Setting yields:
Variation Penalty Bound
For the variation term, note that:
Using the smoothness assumption and the fact that :
Combine Terms
The total dynamic regret is:
Using and simplifying constants gives the stated bound. ∎
C.1.3 Average Performance Guarantee
Theorem 4 (Average Performance).
Let be the average of all iterates. Then:
where .
Proof.
By Jensen’s inequality and convexity of the loss functions:
The first term is bounded by Theorem 3. For the second term, note that and is bounded, giving the stated result. ∎
C.1.4 Justification of Design Choices
Design choice justification: Our algorithm uses , but why linear? We could use exponential, polynomial, or other schedules. The next result shows linear scheduling is actually optimal among reasonable choices.
Proposition 5 (Linear Schedule Optimality).
Among all scheduling functions with , , and bounded total variation , the linear schedule minimizes the variation penalty term in Theorem 3.
Proof.
The variation penalty depends on . For fixed endpoints and bounded total variation, this sum is minimized when increases as slowly as possible initially, which corresponds to the linear schedule. ∎
The convergence analysis above shows our algorithm works, but doesn’t explain why forgetting helps. We now take a domain adaptation perspective: fine-tuning can be viewed as adapting from the pre-training distribution to a target task distribution. Token-level forgetting helps by removing harmful distribution shift effects while preserving beneficial patterns.
C.2 Influence function for Domain Adaptation
We establish rigorous theoretical foundations for our forgetting-based fine-tuning approach in large language models (LLMs). Our analysis extends the distribution shift decomposition framework to understand how proportion-based token selection and selective forgetting improve domain adaptation performance.
C.2.1 Problem Formalization and Setup
Definition 6 (LLM Domain Adaptation Setup).
The goal is to minimize the target risk while preserving performance on the pre-training distribution , where we consider: 1) Pre-training distribution over representing the massive corpus used for initial training, 2) Target domain distribution over representing our specific task data, 3) Model where is the -dimensional probability simplex, and 4) Cross-entropy loss where denotes the probability assigned to token .
Connecting to our method: This formalization captures exactly what happens in LLM fine-tuning. The pre-training distribution represents the massive corpus used for initial training, while represents our specific task data. Our goal is minimizing (good performance on the target task) while not degrading too much (preserving general capabilities). Token-level forgetting helps achieve this balance.
Assumption 7 (Regularity and Smoothness Conditions).
We impose the following regularity conditions:
1.
Lipschitz Loss: For any , and parameters :
2.
Bounded Gradients: For all and :
3.
Bounded Loss: For all and :
4.
Smooth Loss: The loss function is twice differentiable with respect to , and the Hessian satisfies:
Assumption 8 (Distributional Assumptions).
We assume the following about the distributions and :
1.
Support Overlap: (no completely novel tokens in target domain)
2.
Bounded Density Ratios: For all :
for some constant
3.
Finite Samples: We have i.i.d. samples from for training
C.2.2 Extension of Distribution Shift Decomposition to Token-Level Forgetting
We extend the DISDE (Distribution Shift DEcomposition) framework of [77] to understand how token-level interventions affect overall performance in LLM fine-tuning.
Definition 9 (Shared Distribution).
Following [77], given distributions and over , define the shared distribution with density:
This corresponds to the harmonic mean of the two densities, emphasizing regions of high overlap between pre-training and target distributions.
Theorem 10 (Token-Level Extension of DISDE).
Let be the shared distribution from Definition 9, and define the conditional risk for . The standard DISDE decomposition gives:
For proportion-based token forgetting with retention probability , let denote the set of top tokens and the remaining tokens. Selective forgetting modifies these terms as:
Proof.
The standard DISDE decomposition follows from [77]. For the forgetting modification, proportion-based selection with parameter effectively creates a filtered training distribution where only the top proportion of tokens (by influence score) receive gradient updates. This corresponds to indicator weighting in each term of the decomposition, representing the fact that tokens in are selectively forgotten during training. ∎
Remark 11 (Interpretation for Proportion-Based Forgetting).
Theorem 10 shows that proportion-based forgetting affects each component of distribution shift in three distinct ways.
X shift reduction (P → S): Forgetting low-quality tokens reduces harmful frequency shifts from pre-training to shared distribution. When we remove tokens that the model struggles to learn from, we prevent the model from being misled by spurious patterns that appear frequently in pre-training but are not actually predictive. This selective removal helps the model focus on the more reliable patterns that generalize well to the target domain.
Y|X shift preservation: Retaining high-quality tokens preserves beneficial relationship patterns where and agree. The tokens we choose to keep are those where the relationship between input and output is consistent and learnable across both distributions. By maintaining these tokens, we ensure that the model continues to benefit from the knowledge it can reliably transfer from pre-training to the target task.
X shift control (S → Q): Maintaining top tokens ensures adequate coverage of target-specific patterns. The proportion-based selection guarantees that we retain enough tokens to capture the essential patterns needed for good performance on the target distribution, while still filtering out the most harmful examples that could degrade generalization.
The key insight is that proportion controls the trade-off between preserving beneficial patterns and eliminating harmful ones, providing a principled way to balance these competing objectives in domain adaptation.
C.2.3 Influence Function Theory for Token Quality Assessment
We formalize our token quality assessment using influence functions, providing theoretical justification for the proportion-based selection scheme.
Our algorithm needs to distinguish “good” tokens from “bad” ones, but what makes a token good? Intuitively, good tokens are those where the model learns useful, generalizable patterns. We formalize this using influence functions: tokens where the loss decreases significantly from base model to reference model represent learnable patterns.
Definition 12 (Token Influence Score).
For a training dataset , base model , and reference model , define the influence score of token as:
This measures how much more (or less) predictable the token becomes after training, indicating whether it represents a learnable pattern.
Assumption 13 (Influence Function Regularity).
We assume:
1.
Score Boundedness: for some constant
2.
Noise Tolerance: Influence scores are robust to small perturbations in
3.
Ranking Stability: The relative ordering of tokens by influence score is stable across similar model checkpoints
Proposition 14 (Optimal Proportion-Based Token Classification).
Proof.
The proportion-based approach is optimal for several complementary reasons: 1) Adaptive to Distribution since unlike fixed thresholds, proportion-based selection automatically adapts to the score distribution across different datasets and domains, 2) Robust to Scale as there is no need to tune absolute thresholds across different datasets or model scales, which would require extensive hyperparameter search, 3) Guaranteed Coverage because it always retains exactly proportion of training data, ensuring stable training and preventing degenerate cases where too few or too many tokens are selected, and 4) Ranking Optimality where under ranking stability (Assumption 13), selecting top- tokens maximizes the expected benefit from retained tokens while minimizing the harm from discarded tokens. ∎
Our method requires choosing the proportion of tokens to keep. Too high () means keeping noisy tokens; too low () means discarding useful data. This theorem characterizes the optimal choice as balancing target domain performance against source domain degradation.
Theorem 15 (Optimal Retention Proportion).
The optimal retention proportion satisfies:
where:
and denotes parameters trained with proportion , controls source preservation, and typically for LLM fine-tuning tasks.
Proof.
The optimal balances two competing objectives: minimizing target domain loss while preserving source domain performance. The first term encourages selecting only the most beneficial tokens, while the second term (weighted by ) prevents catastrophic forgetting of source capabilities. The typical range emerges from empirical observations that most datasets contain 20-40% of harmful or uninformative tokens. ∎
C.2.4 Generalization Bounds for Proportion-Based Forgetting
Our main theoretical result establishes that proportion-based forgetting provably improves generalization in domain adaptation.
This theorem provides a definitive answer by bounding the target domain risk. The bound reveals how proportion-based forgetting affects generalization through four distinct channels: standard target domain uncertainty, bias from forgetting, transfer error between domains, and selection variance.
Theorem 16 (Main Generalization Bound).
Understanding the bound: Each term in this bound has a clear interpretation. The first term is unavoidable (limited target data), the second shows forgetting bias is controlled by token influence scores, the third captures domain transfer difficulty, and the fourth reflects selection uncertainty. Importantly, the second term can be negative when we forget truly harmful tokens, actually improving the bound.
Proof Sketch.
The proof decomposes the generalization error into four components:
Standard generalization on target domain.
Forgetting-induced bias. The bias from forgetting proportion of tokens:
Transfer error.
Selection variance. The variance from proportion-based selection: ∎
Corollary 17 (Optimal Rate with Proportion-Based Forgetting).
Under the conditions of Theorem 16, proportion-based forgetting achieves:
where represents oracle parameters with perfect domain adaptation. This matches the minimax optimal rate for supervised domain adaptation.
Theorem 18 (Proportion-Specific Bound Analysis).
The generalization bound in Theorem 16 is minimized when:
where is the average absolute influence of negative tokens. This provides theoretical guidance for selecting the retention proportion.
C.2.5 Connection to Distribution Shift Analysis
Theorem 19 (Proportion-Based DISDE Analysis).
Under proportion-based forgetting with parameter , the three components of distribution shift are affected as:
| X shift reduction: | (41) | |||
| Y|X shift preservation: | (42) | |||
| Novelty shift control: | (43) |
where denotes the -th component without forgetting. This shows that proportion-based forgetting provides principled control over each type of distribution shift.